A seller allows a discount of 5% on watch.
if it discounts at 7% he earns Rs.15 less in profit. What is the MP;

Answer by Theo(13342)

profit is equal to revenue minus cost.

sale price = retail price minus discount.

at 5% discount, the sale price is equal to retail price minus 5% of retail price.

at 7% discount, the sale price is equal to retail price minus 7% of retail price.

when he sells it at 5% discount, he makes 15 more than when he sells it at 7% discount.

this means that sale price at 5% discount minus sale price at 7% discount = 15.

the difference between the sale price at 5% discount and the sale price at 7% discount is equal to 2% of the retail price.

this means that 2% of the retail price is equal to 15.

this means that the retail price is 15 / 2% = 15 / .02 = 750.

the sale price at 5% discount = 750 - .05 * 750 = 712.5
the sale price at 7% discount = 750 - .07 * 750 = 697.5

in algebraic terms, you would solve it something like this:

r = retail price
s1 = sale price at 5% discount
s2 = sale price at 7% discount
d1 = .05r
d2 = .07r
p1 = profit at 5% discount
p2 = profit at 7% discount
c = cost

s1 = r-d1
s2 = r-d2

since d1 = .05r and d2 = .07r, this leads to:

s1 = r - .05r
s2 = r - .07r

profit at 5% discount = sale price at 5% discount - cost
profit at 7% discount = sale price at 7% discount - cost

since c is the cost, this leads to:

p1 = s1 - c
p2 = s1 - c

since s1 = r-.05r and s2 = r-.07r, this leads to:

p1 = r-.05r-c
p2 = r-.07r-c

since profit at 5% discount is 15 more than profit at 7% discount, this leads to:

p1-p2=15

since p1 = r-.05r-c and p2 = r-.07r-c, this leads to:

(r-.05r-c) - (r-.07r-c) = 15

remove parentheses to get:

r-.05r-c-r+.07r+c = 15

the r and the c cancel out and you are left with:

-.05r+.07r = 15

combine like terms to get:

.02r = 15

solve for r to get r = 15/.02 = 750.

that's the retail price.

the sale price at 5% discount = 750 - .05*750 = 712.5
the sale price at 7% discount = 750 - .07*750 = 697.5

i believe the sale price is the same as what you call the market price.

your solution depends on what market price they are looking for.